OpenStudy (calculusxy):

How many two digit whole numbers are increased by 18 when their digits are reversed? @phi

OpenStudy (kainui):

I look at this and think, what can I easily eliminate? So I ask you, what's your best guess? You have to start somewhere, so take some guesses and I'll help point you in the right direction.

OpenStudy (calculusxy):

The part that I find confusing is when it says about the number being increased by 18, but then asks about how many of them of are increased?

OpenStudy (kainui):

So all this is saying is suppose you have a number like 12 and reverse the digits, you get 21. That's an increase 9.

OpenStudy (calculusxy):

Can you explain to me more about how is it increased by 9?

OpenStudy (kainui):

12+9=21

OpenStudy (calculusxy):

Ok. And then what do i do?

OpenStudy (phi):

I would write an equation. if you have the "number" ab, it has the value 10a+b it goes up by 18 if you reverse the digits 10a+b+18 = 10b + a

OpenStudy (kainui):

Then you use this nice little example I've laid out for you to think about it and figure it out.

OpenStudy (calculusxy):

@phi Why is it next to 10 (meaning ab)?

OpenStudy (phi):

the number 23 has the value 2*10 + 3

OpenStudy (calculusxy):

Ok

OpenStudy (phi):

simplify 10a+b+18 = 10b + a to a= b-2 now list all digits that b can be, and find the corresponding a

OpenStudy (calculusxy):

Would it be 9a+18=9b?

OpenStudy (anonymous):

ab - ba= 10a + b - 10b - a = 18 --> 9(a-b) = 18 --> a-b=2. In whole numbers, for example a, with 1 digit, we have totally 10 possible cases (0,1,2,3 ---9) But a and b cannot be equal to 0, then only 9 possible cases left. a - b = 2 which means a > 2 also, therefore a only takes 3 --> 9 (7 possible values) For 7 possible values of a, we have 7 possible values of b. Therefore, totally the answer is we have 7 possible two-digit whole numbers (couples) which reverse for difference of 18 They are: 31 and 13 ; 42 and 24; 53 and 35; 64 and 46; 75 and 57; 86 and 68; 97 and 79

OpenStudy (phi):

yes, now divide by 9 to get a+2 = b

OpenStudy (phi):

I guess we have to leave out 02 (+18= 20) because 02 is not a two digit number.

OpenStudy (phi):

but you can just use b= 0... 9 and toss out the invalid ones. example a+2 = 0 -> a= -2 not allowed...

OpenStudy (phi):

or start with a=0 ... 9 and find b. Again, toss out invalid numbers a=0--> b=2 but 02 is not allowed because it is not a two digit number.

OpenStudy (calculusxy):

I don't get about what I should do with a+2=b. Can you example to me this more?

OpenStudy (phi):

let a=0 ... 9 and find b. using a+2 = b

OpenStudy (calculusxy):

Like a=1 + 2 and then b=3. 1+2=3

OpenStudy (phi):

*** Like a=1 + 2 and then b=3. 1+2=3 *** yes. that means your number is 13 if we add 18 to it we should get 31... and we do. so that is one of the solutions

OpenStudy (calculusxy):

Oh! Can you help me approach all the solutions?while I do the work though. I don't need answers yet :)

OpenStudy (phi):

just keep going and writing down what you get... (the answers will be "ab" concatenated together)

OpenStudy (calculusxy):

Ok thanks

OpenStudy (phi):

start with b= a+2 make a list of all digits a can be 0,1,2... up to 9 for each a digit, write the "b" digit next to it (b is a+2) cross off any combination that does not make sense.

OpenStudy (kainui):

@phi you listed "0" as being a possible digit, but remember that's not possible if these are two digit numbers.

OpenStudy (phi):

yes, we will cross off 0... but in the beginning I would list all possibilities. First do the mechanical part... then do the "thinking" part

OpenStudy (kainui):

Seems like once you've narrowed it down to a+2=b you've really got a nice clean shot by thinking rather than going through any kind of hardship. a and b are both single digit numbers, so we can finish up by making these extra couple of equations: 0<a<10 0<b<10 And this combined with a+2=b will give you all the information you need.

OpenStudy (phi):

good point... I was assuming it was self-evident that a and b are restricted to single digits but it is always good to make that explicit.

OpenStudy (calculusxy):

Ok so I came up to like 32+18=50. That means that that won't work right?

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

How did you get 32 ?

OpenStudy (calculusxy):

3(10)+2=32

OpenStudy (phi):

make this table a b= a+2 ab 1 3 13 2 4 24 and so on...

OpenStudy (calculusxy):

Thank you soooooo much. I got my answer as 7. And that was correct!

OpenStudy (anonymous):

14 A Mathematica solution is attached.